# Examining Input Modalities and Visual Feedback Designs in Mobile Expressive Writing

### v1

@@ 7.3. Research Opportunities in Mobile Expressive Writing
The affordances of modern smartphones are different from those of pen and paper — the original tools of expressive writing. Prior work has shown that expressive writing helps release repressed emotions, reorganize feelings and calm down (Pennebaker and Chung, 2007). Yet, most of past work has naturally been conducted with tools other than the relatively modern ones, such as smartphones. Our results with digital smartphone-powered expressive writing thus contribute to new empirical understanding of how interface factors of mobile technology can influence expressive writing effect and experience.
The qualitative findings in the second study suggest that different visual designs offer different experiences for the users. People felt a different level of agency over relieving their stress; participants felt as if they are observers in the Passive condition while they felt like they are actively destroying their stress in the Interactive condition. Existing literature suggests that personality (e.g., extrovert vs. introvert) could influence preferences and use scenarios of different visual feedback (Carver and Connor-Smith, 2010). A more extensive study considering these and other similar factors will lead into a deeper understanding of appropriate interface designs for different users and use cases.
The latest speech-based mobile input technologies can perform real-time recognition and transcription (Li et al., 2022). Engaging in a sensitive task, such as expressive writing, calls for exploring optimal input methods. Our results between text and speech based input showed clear differences, as explored in Section 4.1. Related to this, Luo demonstrated a combination of keyboard-based and speech-based text entry for food journaling (Luo, 2022), and such mixture may obtain a more optimal experience with respect to usability and stimulation of self-reflection. Examining the division of input tasks between keyboard-based and speech-based text entry for mobile expressive writing is yet another an interesting future research direction.
Overall, our results suggest that smartphones can facilitate effective expressive writing. Given their unique features, they can facilitate many types of rich interactions that are not available with other writing platforms. Auditory or haptic feedback (Takami et al., 2022) may be a possible direction to explore with expressive writing solutions For instance, Costa et al. (Costa et al., 2016) showed how haptic feedback of heart rate could help users regulate their anxiety. Future systems could include deep breathing and mindfulness activities with emotional expression and examine how they could enhance cathartic effects and overall stress relief in conjunction with expressive writing tasks. Such multimodal, multi-approach experiences are fairly trivial to program with modern smartphones coupled with off-the-shelf wearable sensors, such as heart rate monitors. Multimodal interface and feedback designs for stress relief may thus lead to a new interdisciplinary research opportunity for HCI and psychology.
@@ 7.4. Limitations
We acknowledge limitations in our work. We conducted our studies in Japan, and they required proficient Japanese communication skills. Accordingly, all participants were Japanese, constituting a homogeneous sample pool with respect to cultural backgrounds. This cultural specificity may have influenced the effectiveness of speech input as talking does not strongly relate to thinking in some Asian cultures, unlike in Western contexts (Kim, 2002). Future studies should examine how speech input could offer different effects and user experience in other cultural contexts. This work would serve as a reference for such studies to uncover the influence of cultural differences on mobile expressive writing.
Another homogeneity in our participants was that none of our participants regularly used speech input with their smartphones, which might have affected the preference results. However, our qualitative results uncover reasons of preferences that go beyond the established familiarity to software keyboards, and thus confirm the advantage of keyboard-based text entry for mobile expressive writing. Our results encourage researchers to further explore input modality effects in different cultural backgrounds and user populations.
The studies reported in this article are also rather short, compared to some of the seminal work in expressive writing in the field of psychology. Yet, during the time span our participants used the developed solutions, we were able to collect sufficiently data for quantitative and qualitative account of how the input modalities and visual designs differed. We argue these studies adequately contribute to the HCI literature on exploring smartphones as an expressive writing tool for cathartic effect, even if not so much on the already well-established method of expressive writing itself.
@@ 8. Conclusion
Expressing negative emotions is known to be an effective stress-coping mechanism, but challenges with respect to interface designs emerge when the practice is transitioned to technology-supported mental well-being systems. This work investigated how input modalities for emotional expression and visual feedback designs for post-expression interventions could affect the perceived cathartic effects and user experience. Our study highlighted the benefits and stronger preference for keyboard-based text entry over input methods using voice messages, and also revealed how different visual feedback designs can benefit users to accommodate different scenarios of stress-related emotional expression. Our findings also suggest important design considerations and future research directions in mobile expressive writing, and further encourage interdisciplinary research between HCI and self-help approach in psychology.
@@ Acknowledgements.
We are grateful for the support received during the preparation of this paper. In particular, we would like to thank Ginshi Shimojima, Chi-lan Yang, Zefan Sramek, and Shitao Fang for their valuable feedback. We also extend our appreciation to the participants in our user study for their involvement. This research is part of the results of Value Exchange Engineering, a joint research project between R4D, Mercari, Inc., and the RIISE. This project was supported by JSPS Bilateral (Grant Number: 120232701).References
@@
- Abd-Alrazaq et al. (2021) Alaa A Abd-Alrazaq, Mohannad Alajlani, Nashva Ali, Kerstin Denecke, Bridgette M Bewick, and Mowafa Househ. 2021. Perceptions and opinions of patients about mental health chatbots: scoping review. Journal of medical Internet research 23, 1 (2021), e17828.
- Ahmed et al. (2022) Arfan Ahmed, Sarah Aziz, Mohamed Khalifa, Uzair Shah, Asma Hassan, Alaa Abd-Alrazaq, and Mowafa Househ. 2022. Thematic analysis on user reviews for depression and anxiety chatbot apps: Machine learning approach. JMIR Formative Research 6, 3 (2022), e27654.
@@
- Hammen (2005) Constance Hammen. 2005. Stress and depression. Annu. Rev. Clin. Psychol. 1 (2005), 293–319.
@@
- Mushtaq et al. (2018) Mamoona Mushtaq et al. 2018. Stress, headache, high blood pressure and self-esteem in working women. Pakistan Heart Journal 51, 2 (2018).
@@
- Sas et al. (2016) Corina Sas, Steve Whittaker, and John Zimmerman. 2016. Design for rituals of letting go: An embodiment perspective on disposal practices informed by grief therapy. ACM Transactions on Computer-Human Interaction (TOCHI) 23, 4 (2016), 1–37.
@@
- Veit and Ware (1983) Clairice T Veit and John E Ware. 1983. The structure of psychological distress and well-being in general populations. Journal of consulting and clinical psychology 51, 5 (1983), 730.
- Walther (1993) Joseph B Walther. 1993. Impression development in computer-mediated interaction. Western Journal of Communication (includes Communication Reports) 57, 4 (1993), 381–398.
@@
- Xu et al. (2012) Alison Jing Xu, Rami Zwick, and Norbert Schwarz. 2012. Washing away your (good or bad) luck: Physical cleansing affects risk-taking behavior. Journal of Experimental Psychology: General 141, 1 (2012), 26.
@@
- Zhong and Liljenquist (2006) Chen-Bo Zhong and Katie Liljenquist. 2006. Washing away your sins: Threatened morality and physical cleansing. Science 313, 5792 (2006), 1451–1452.
@@ Appendix A Questionnaires Used in Study 1
The original questionnaire was in Japanese, and we translated it into English for the report in this paper.
@@ Appendix B Questionnaires Used in Study 2
The original questionnaire was in Japanese, and we translated it into English for the report in this paper.

### v3

@@ 7.3. Research Opportunities in Mobile Expressive Writing
Our second study indicates that visual design affects user experiences differently. Participants felt like observers in the Passive condition and like active stress relievers in the Interactive condition. Literature suggests personality (e.g., extrovert vs. introvert) influences preferences for visual feedback (Carver and Connor-Smith, 2010). A more extensive study on these factors can further deepen the understanding of mobile interface designs for expressive writing. Further, when our participants received text-based feedback after disclosure, they felt like someone heard their narratives. Here, future research could investigate how to design feedback so that users focus more on the disclosed emotions themselves rather than the feedback and its effects, which could lead to useful cognitive restructuring.
Speech-based mobile input can provide real-time recognition and transcription (Li et al., 2022). As our study identified clear differences between text and speech inputs (Section 4.1), sensitive tasks like expressive writing require careful design considerations on input methods. Luo showed that combining keyboard and speech input for food journaling (Luo, 2022) could enhance usability and self-reflection. Future research could explore dividing tasks between keyboard and speech input for mobile expressive writing.
Smartphones can facilitate many types of rich interactions that are not available with other writing platforms. Auditory or haptic feedback (Takami et al., 2022) may be a possible direction to explore with expressive writing solutions For instance, Costa et al. (Costa et al., 2016) showed how haptic feedback of heart rate could help users regulate their anxiety. Future systems could include deep breathing and mindfulness activities with emotional expression and examine how they could enhance cathartic effects and overall stress relief in conjunction with expressive writing tasks. Multimodal interface and feedback designs for stress relief may thus lead to a new interdisciplinary research opportunity for HCI and psychology. Moreover, smartphones are excellently equipped for longitudinal data collection studies. While our focus here is on comparing the conditions (designs), a longitudinal study with a similar journaling application focusing on temporal changes would be interesting. Here, issues such as how participants perceptions evolve as a function of study progression, or whether their preferences shift over time would reveal important findings about the effects of novelty in similar deployments.
@@ 7.4. Limitations
We acknowledge several limitations in our work. Our study was conducted in Japan with a culturally homogeneous sample of native Japanese speakers. This might have influenced speech input effectiveness, as verbalizing thoughts relates differently to cognitive processing in some Asian cultures compared to the West (Kim, 2002). Future research should consider diverse cultural backgrounds. None of the participants frequently used speech input, which may have affected their preferences. Further, our qualitative findings indicate that keyboard input advantages extend beyond mere familiarity. We adapted questionnaires used in previous work to our studies, and we used the same questionnaires across the conditions to minimize the effect of question designs. Future work may further validate our results by employing different framing and styles of questions. Despite the limitations, our work contributes two successful deployment studies on expressive writing in authentic settings, which advance the field of mobile HCI for stress management.
@@ 8. Conclusion
Expressing negative emotions is known to be an effective stress-coping mechanism, but challenges with respect to interface designs emerge when the practice is transitioned to technology-supported mental well-being systems. This work investigated how input modalities for emotional expression and visual feedback designs for post-expression interventions could affect the perceived cathartic effects and user experience. Our study highlighted the benefits and stronger preference for keyboard-based text entry over input methods using voice messages. It also revealed how different visual feedback designs can benefit users to accommodate different scenarios of stress-related emotional expression. Our findings also suggest important design considerations and future research directions in mobile expressive writing, and further encourage interdisciplinary research between HCI and self-help approach in psychology.
@@ Acknowledgements.
We are grateful for the support received during the preparation of this paper. In particular, we would like to thank Ginshi Shimojima, Chi-lan Yang, Zefan Sramek, and Shitao Fang for their valuable feedback. We also extend our appreciation to the participants in our user study for their involvement. This research is part of the results of Value Exchange Engineering, a joint research project between R4D, Mercari, Inc., and the RIISE. This project was also partly supported by JSPS Bilateral (Grant Number: 120232701), Research Council of Finland (Grant Numbers 349637 and 353790), and JST ASPIRE for Top Scientists (Grant Number: JPMJAP2405).References
@@
- Alhasani et al. (2022) Mona Alhasani, Dinesh Mulchandani, Oladapo Oyebode, Nilufar Baghaei, and Rita Orji. 2022. A systematic and comparative review of behavior change strategies in stress management apps: opportunities for improvement. Frontiers in Public Health 10 (2022), 777567.
- Axsom et al. (1987) Danny Axsom, Suzanne Yates, and Shelly Chaiken. 1987. Audience response as a heuristic cue in persuasion. Journal of personality and social psychology 53, 1 (1987), 30.
@@
- Bond and Pennebaker (2012) Mark Bond and James W Pennebaker. 2012. Automated computer-based feedback in expressive writing. Computers in Human Behavior 28, 3 (2012), 1014–1018.
@@
- Burton and King (2004) Chad M Burton and Laura A King. 2004. The health benefits of writing about intensely positive experiences. Journal of research in personality 38, 2 (2004), 150–163.
- Burton and King (2008) Chad M Burton and Laura A King. 2008. Effects of (very) brief writing on health: The two-minute miracle. British Journal of Health Psychology 13, 1 (2008), 9–14.
@@
- Cherubini et al. (2009) Mauro Cherubini, Xavier Anguera, Nuria Oliver, and Rodrigo De Oliveira. 2009. Text versus speech: a comparison of tagging input modalities for camera phones. In Proceedings of the 11th International Conference on Human-Computer Interaction with Mobile Devices and Services. 1–10.
- Chung and Pennebaker (2008) Cindy K Chung and James W Pennebaker. 2008. Variations in the spacing of expressive writing sessions. British Journal of Health Psychology 13, 1 (2008), 15–21.
@@
- Daiute (1986) Colette Daiute. 1986. Physical and cognitive factors in revising: Insights from studies with computers. Research in the Teaching of English 20, 2 (1986), 141–159.
@@
- Fitzpatrick et al. (2017) Kathleen Kara Fitzpatrick, Alison Darcy, and Molly Vierhile. 2017. Delivering cognitive behavior therapy to young adults with symptoms of depression and anxiety using a fully automated conversational agent (Woebot): a randomized controlled trial. JMIR mental health 4, 2 (2017), e7785.
- Flower and Hayes (1981) Linda Flower and John R Hayes. 1981. A cognitive process theory of writing. College Composition & Communication 32, 4 (1981), 365–387.
@@
- Gersten and Baker (2001) Russell Gersten and Scott Baker. 2001. Teaching expressive writing to students with learning disabilities: A meta-analysis. The elementary school journal 101, 3 (2001), 251–272.
- Greenberg et al. (1996) MA Greenberg, AA Stone, and CB Wortman. 1996. Health and psychological effects of emotional disclosure: A test of the inhibition-confrontation approach. Journal of Personality and Social Psychology 71, 3 (1996), 588–602.
@@
- Han et al. (2022) Yunjo Han, Hyemin Lee, Kobiljon E Toshnazarov, Youngtae Noh, and Uichin Lee. 2022. Stressbal: personalized just-in-time stress intervention with wearable and phone sensing. In Adjunct Proceedings of the 2022 ACM International Joint Conference on Pervasive and Ubiquitous Computing and the 2022 ACM International Symposium on Wearable Computers. 41–43.
@@
- Harrist et al. (2007) Steve Harrist, Barbara L Carlozzi, Arthur R McGovern, and Amanda W Harrist. 2007. Benefits of expressive writing and expressive talking about life goals. Journal of Research in Personality 41, 4 (2007), 923–930.
@@
- Hatvany et al. (2018) Thomas Hatvany, Edward Burkley, and Jessica Curtis. 2018. Becoming part of me: Examining when objects, thoughts, goals, and people become fused with the self-concept. Social and Personality Psychology Compass 12, 1 (2018), e12369.
- Higgs and Joseph (1971) WJ Higgs and Karen B Joseph. 1971. Effects of real and anticipated audiences on verbal learning and reproduction. The Journal of Social Psychology 85, 1 (1971), 41–49.
@@
- Juarascio et al. (2021) Adrienne Juarascio, Paakhi Srivastava, Emily Presseller, Kelsey Clark, Stephanie Manasse, Evan Forman, et al. 2021. A clinician-controlled just-in-time adaptive intervention system (CBT+) designed to promote acquisition and utilization of cognitive behavioral therapy skills in bulimia nervosa: Development and preliminary evaluation study. JMIR Formative Research 5, 5 (2021), e18261.
- Kanaya and Kawai (2024) Yuta Kanaya and Nobuyuki Kawai. 2024. Anger is eliminated with the disposal of a paper written because of provocation. Scientific Reports 14, 1 (2024), 7490.
- Keltner and Bonanno (1997) Dacher Keltner and George A Bonanno. 1997. A study of laughter and dissociation: distinct correlates of laughter and smiling during bereavement. Journal of personality and social psychology 73, 4 (1997), 687.
@@
- King and Miner (2000) Laura A King and Kathi N Miner. 2000. Writing about the perceived benefits of traumatic events: Implications for physical health. Personality and social psychology bulletin 26, 2 (2000), 220–230.
- Kliewer et al. (2011) Wendy Kliewer, Stephen J Lepore, Albert D Farrell, Kevin W Allison, Aleta L Meyer, Terri N Sullivan, and Anne Y Greene. 2011. A school-based expressive writing intervention for at-risk urban adolescents’ aggressive behavior and emotional lability. Journal of Clinical Child & Adolescent Psychology 40, 5 (2011), 693–705.
@@
- Lai and Yankelovich (2006) Jennifer Lai and Nicole Yankelovich. 2006. Speech interface design. (2006).
- Langer et al. (2012) Shelby L Langer, Thomas H Kelly, Barry E Storer, Suzanne P Hall, Heather G Lucas, and Karen L Syrjala. 2012. Expressive talking among caregivers of hematopoietic stem cell transplant survivors: acceptability and concurrent subjective, objective, and physiologic indicators of emotion. Journal of psychosocial oncology 30, 3 (2012), 294–315.
@@
- Lord et al. (1987) Charles G Lord, Delia S Saenz, and Debra K Godfrey. 1987. Effects of perceived scrutiny on participant memory for social interactions. Journal of Experimental Social Psychology 23, 6 (1987), 498–517.
@@
- Luo et al. (2022) Yuhan Luo, Bongshin Lee, Young-Ho Kim, and Eun Kyoung Choe. 2022. NoteWordy: Investigating touch and speech input on smartphones for personal data capture. Proceedings of the ACM on Human-Computer Interaction 6, ISS (2022), 568–591.
- Lyubomirsky et al. (2006) Sonja Lyubomirsky, Lorie Sousa, and Rene Dickerhoof. 2006. The costs and benefits of writing, talking, and thinking about life’s triumphs and defeats. Journal of personality and social psychology 90, 4 (2006), 692.
- MacArthur et al. (1996) Charles A MacArthur, Shirley S Schwartz, Steve Graham, Dawn Molloy, et al. 1996. Integration of strategy instruction into a whole language classroom: A case study. Learning Disabilities Research & Practice (1996).
@@
- Moore (2016) James W Moore. 2016. What is the sense of agency and why does it matter? Frontiers in psychology 7 (2016), 1272.
@@
- Nahum-Shani et al. (2018) Inbal Nahum-Shani, Shawna N Smith, Bonnie J Spring, Linda M Collins, Katie Witkiewitz, Ambuj Tewari, and Susan A Murphy. 2018. Just-in-time adaptive interventions (JITAIs) in mobile health: key components and design principles for ongoing health behavior support. Annals of Behavioral Medicine 52, 6 (2018), 446–462.
- Newton (2015) Kevin Michael Newton. 2015. Emotional Disclosure: Does Writing Differ from Speaking? (2015).
- Niles et al. (2016) Andrea N Niles, Kate E Byrne Haltom, Matthew D Lieberman, Christopher Hur, and Annette L Stanton. 2016. Writing content predicts benefit from written expressive disclosure: Evidence for repeated exposure and self-affirmation. Cognition and Emotion 30, 2 (2016), 258–274.
@@
- O’Leary et al. (2018) Kathleen O’Leary, Stephen M Schueller, Jacob O Wobbrock, and Wanda Pratt. 2018. “Suddenly, we got to become therapists for each other” Designing Peer Support Chats for Mental Health. In Proceedings of the 2018 CHI conference on human factors in computing systems. 1–14.
@@
- Presseller et al. (2022) Emily K Presseller, Megan L Wilkinson, Claire Trainor, Elizabeth W Lampe, and Adrienne S Juarascio. 2022. Self-regulation deficits moderate treatment outcomes in a clinical trial evaluating just-in-time adaptive interventions as an augmentation to cognitive-behavioral therapy for bulimia-spectrum eating disorders. International Journal of Eating Disorders 55, 5 (2022), 709–716.
@@
- Reynolds et al. (2000) Martina Reynolds, Chris R Brewin, and Matthew Saxton. 2000. Emotional disclosure in school children. The Journal of Child Psychology and Psychiatry and Allied Disciplines 41, 2 (2000), 151–159.
@@
- Rothbart (1973) Mary K Rothbart. 1973. Laughter in young children. Psychological bulletin 80, 3 (1973), 247.
@@
- Sawyer et al. (1992) Richard J Sawyer, Steve Graham, and Karen R Harris. 1992. Direct teaching, strategy instruction, and strategy instruction with explicit self-regulation: Effects on the composition skills and self-efficacy of students with learning disabilities. Journal of educational psychology 84, 3 (1992), 340.
- Schalkwyk et al. (2010) Johan Schalkwyk, Doug Beeferman, Françoise Beaufays, Bill Byrne, Ciprian Chelba, Mike Cohen, Maryam Kamvar, and Brian Strope. 2010. “Your word is my command”: Google search by voice: A case study. In Advances in Speech Recognition: Mobile Environments, Call Centers and Clinics. Springer, 61–90.
@@
- Smyth and Pennebaker (2008) Joshua M Smyth and James W Pennebaker. 2008. Exploring the boundary conditions of expressive writing: In search of the right recipe. British journal of health psychology 13, 1 (2008).
@@
- Soliday et al. (2004) Elizabeth Soliday, John P Garofalo, and Debra Rogers. 2004. Expressive writing intervention for adolescents’ somatic symptoms and mood. Journal of Clinical Child and Adolescent Psychology 33, 4 (2004), 792–801.
@@
- van Genugten et al. (2025) Claire R van Genugten, Melissa SY Thong, Wouter Van Ballegooijen, Annet M Kleiboer, Donna Spruijt-Metz, Arnout C Smit, Mirjam AG Sprangers, Yannik Terhorst, and Heleen Riper. 2025. Beyond the current state of just-in-time adaptive interventions in mental health: a qualitative systematic review. Frontiers in Digital Health 7 (2025), 1460167.
- Wakschlag et al. (2021) Lauren S Wakschlag, Darius Tandon, Sheila Krogh-Jespersen, Amelie Petitclerc, Ashley Nielsen, Rhoozbeh Ghaffari, Leena Mithal, Michael Bass, Erin Ward, Jonathan Berken, et al. 2021. Moving the dial on prenatal stress mechanisms of neurodevelopmental vulnerability to mental health problems: A personalized prevention proof of concept. Developmental psychobiology 63, 4 (2021), 622–640.
@@
- Wong et al. (1996) Bernice YL Wong, Deborah L Butler, Sheryl A Ficzere, and Sonia Kuperis. 1996. Teaching low achievers and students with learning disabilities to plan, write, and revise opinion essays. Journal of learning disabilities 29, 2 (1996), 197–212.
@@
- Zhao et al. (2023) Yiran Zhao, Yujie Tao, Grace Le, Rui Maki, Alexander Adams, Pedro Lopes, and Tanzeem Choudhury. 2023. Affective touch as immediate and passive wearable intervention. Proceedings of the ACM on Interactive, Mobile, Wearable and Ubiquitous Technologies 6, 4 (2023), 1–23.
@@ Appendix A Questionnaires Used in Study 1
The original questionnaire was in Japanese, and we translated it into English for the report in this paper. We referred to our system as “a chatbot” as the interface was integrated into LINE text chat.
@@ Appendix B Questionnaires Used in Study 2
The original questionnaire was in Japanese, and we translated it into English for the report in this paper. We referred to our system as “a chatbot” as the interface was integrated into LINE text chat.